# The torus plateau for the high-dimensional Ising model

### v1

@@ Abstract
We consider the Ising model on a $d$-dimensional discrete torus of volume $r^{d}$, in dimensions $d>4$ and for large $r$, in the vicinity of the infinite-volume critical point $\beta_{c}$. We prove that for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ (with a suitable constant) the susceptibility is bounded above and below by multiples of $r^{d/2}$, and that the two-point function has a “plateau” in the sense that it decays like $|x|^{-(d-2)}$ when $|x|$ is small relative to the volume but for larger $|x|$ it levels off to a constant value of order $r^{-d/2}$. We also prove that at $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ the renormalised coupling constant is nonzero, which implies a non-Gaussian limit for the average spin. The proof relies on near-critical estimates for the infinite-volume two-point function obtained recently by Duminil-Copin and Panis, and builds upon a strategy proposed by Papathanakos. The random current representation of the Ising model plays a central role in our analysis.
@@ 1.1 Introduction
In recent years, the plateau for PBC has been proven to exist for simple random walk in dimensions $d>2$ [40, 15], self-avoiding walk for $d>4$ [40, 30], and percolation for $d>6$ [27, 28]. A general theory of the effect of FBC vs PBC in the setting of the weakly-coupled hierarchical $n$-component $|\varphi|^{4}$ model in dimensions $d\geq 4$ was developed in [33, 37], and we comment further on this below.
For the Ising model in dimensions $d>4$, the absence of a plateau for FBC was proven in [13]. For PBC, incomplete steps were taken in [36] towards proving existence of the plateau. In this paper, we build on the method of [36] to prove (1.8) for PBC in dimensions $d>4$. We also prove the non-vanishing of the renormalised coupling constant under PBC, indicating a non-Gaussian limit for the average spin. This can be contrasted with the Gaussian limit under FBC obtained in [2, 20, 4, 35].
We emphasise that we do not use the lace expansion in any way, and our results do not require a small parameter. The input from the lace expansion used to establish plateau theorems for self-avoiding walk and percolation in [40, 30, 27, 28] is replaced here by the near-critical bounds (1.6)–(1.7).
@@ 1.2 Main results
Let $d>4$. There are constants $c_{i}>0$, depending on $d$ but not on $\beta$, such that
@@ 1.3 Conjectured asymptotic formulas
A precise and detailed account of the effect of boundary conditions on finite-size scaling is given in [33, 37], via a rigorous renormalisation group analysis. The results of [33, 37] are proved for the weakly-coupled $n$-component $|\varphi|^{4}$ model on the $d$-dimensional hierarchical lattice for $d\geq 4$ and for all $n\geq 1$, but, on the basis of universality, the results obtained there for the $1$-component model are conjectured to apply also to all $1$-component spin models in the universality class of the nearest-neighbour Ising model. An antecedent of the conjectures goes back to [12] (see also [43, Chapter 32]). The conjectures present predictions for a specific and universal non-Gaussian limit for the average field, and precise asymptotic amplitudes for the susceptibility, the plateau, and the renormalised coupling constant, throughout a critical window of width $r^{-d/2}$ (with a logarithmic correction for $d=4$) centred at $\beta_{c}$. We define the window scale and the large-field scale, respectively, by
@@
The function $f$ is strictly monotone increasing, and satisfies $f(s)\sim|s|^{-1}$ as $s\to-\infty$ and $f(s)\sim s$ as $s\to\infty$ (see [33, Section 1.5.2] where $f_{1}(s)=f(-s)$). The following conjecture proposes a precise non-Gaussian limit for the average field $S_{r}$, and a precise role for the universal profile $f$ in the critical finite-size scaling of the susceptibility, the two-point function, and the renormalised coupling constant for the Ising model. The conjecture is a special case ($n=1$) of conjectures from [33, 37] for $n$-component models for all $n\geq 0$ ($n=0$ corresponds to the self-avoiding walk). The profile $f$ also arises in the Curie–Weiss model; see [18, Theorem 3.2].
@@
As observed in [33, Section 1.5.3], the right-hand side of (1.20) is strictly monotone increasing, and has limits 00 and $2$ as $s\to-\infty$ and $s\to+\infty$ respectively. Thus, assuming (1.20), the renormalised coupling constant varies monotonically over its possible range $[0,2]$ as the window is traversed. At the infinite-volume critical point, (1.20) gives
@@
such that the statements of Conjecture 1.5 holds as stated when $\beta_{c}$ is replaced by $\beta_{c,r}$. In particular, we conjecture that the plateau phenomenon also occurs for free boundary conditions at the effective critical point.
@@ 1.4 Structure of proof
Let $d\geq 1$, $r\geq 3$, and $x\in{\mathbb{T}_{r}}$. Then $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x).$
@@
The upper bound follows from (1.9), $\tau_{\beta}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ by (1.6), and the fact that $\chi(\beta_{*})\asymp r^{-d/2}$ from (1.5).
@@
Let $d\geq 1$. For all $\beta>0$ and for all $r\geq 3$,
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)|\geq\frac{1}{128d^{2}}\cdot\min\Big{\{}\beta(\chi^{\mathbb{T}_{r}})^{2}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta},\frac{1}{{\sf B}}\Big{(}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta}\Big{)}^{2}\Big{\}}.$ (1.34)
@@ 2.2 Diagrammatic estimates
Let $d\geq 1$. For any $x\in{\mathbb{T}_{r}}$,
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq 2\Big{(}T_{\beta}^{(r)}\star\big{(}T_{\beta}^{(r)}\big{)}^{2}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]\star T_{\beta}^{(r)}\Big{)}(x).$ (2.10)
@@
Since every point in ${\mathbb{Z}^{d}}$ can be written uniquely in the form $\Lambda_{r}+r{\mathbb{Z}^{d}}$, we can write
@@
so the two sums become four sums over $\check{y},\check{z},u,v$. Since $T_{\beta}^{(r)}$ is periodic, we have $T_{\beta}^{(r)}(y-z)=T_{\beta}^{(r)}(\check{y}-\check{z})$ and $T_{\beta}^{(r)}(x-y)=T_{\beta}^{(r)}(\check{x}-\check{y})$, independent of $u$. Hence, the sum over $u$ gives the upper bound
@@
It is interesting to compare the diagrammatic upper bound (2.10) with diagrams in related models. The counterpart of $T_{\beta}^{(r)}-\tau_{\beta}^{\mathbb{T}_{r}}$ is denoted by $\psi-\psi^{\mathbb{T}}$ in [40, 28, 30]. For weakly or strictly self-avoiding walk, the bounds [40, (6.36)] and [30, (4.20)] can be written as
@@
It follows from $\tau_{\beta}(x)\leq\tau_{\beta_{c}}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ and a standard convolution estimate [23, Proposition 1.7] that $(\tau_{\beta}*\tau_{\beta}^{2}*\tau_{\beta})(x)\lesssim|\!|\!|x|\!|\!|^{-(d-4)}$ when $d>4$, so this term can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ with $c_{5}$ large enough, uniformly in $\beta\leq\beta_{c}$. Also, for $\beta\leq\beta_{c}-c_{4}r^{-d/2}$, by (1.5) we have $r^{-d/2}\chi(\beta)\lesssim c_{4}^{-1}$, and we can make this as small as desired by taking $c_{4}$ large enough. With (2.18), this shows that the multiplier of ${\chi(\beta)}/r^{d}$ in (2.17) can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ and $c_{4},c_{5}$ large. In particular, it can be made smaller than $\frac{1}{2}c_{T}$, which completes the proof. ∎
@@
In this section, we introduce a new coupling between Ising models on ${\mathbb{T}_{r}}$ and on large finite subsets of $\mathbb{Z}^{d}$. The coupling uses the random current representation (see [16] for an introduction) and a percolation exploration process. To derive useful bounds, we develop a new switching lemma between currents on $\mathbb{T}_{r}$ and on large finite subsets of $\mathbb{Z}^{d}$, in Lemma 3.4.
@@ 3.1 Random currents and switching lemma
$w_{\beta}(\mathbf{n})=\prod_{xy\in E}\dfrac{\beta^{\mathbf{n}_{xy}}}{\mathbf{n}_{xy}!},\qquad Z^{A}_{\mathcal{G},\beta}=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n}).$ (3.3)
@@
We will use the following minor extension of the switching lemma of [21, 2] (see also [16, Lemma 7]). Given two graphs $\mathcal{G},\mathcal{H}$, two vertex sets $A\subset\mathcal{G}$, $B\subset\mathcal{H}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}\times\Omega_{\mathcal{H}}$, we write
@@
We often drop the subscript $\beta$ when it does not play a role. For two sets $S_{1},S_{2}$, we write $S_{1}\Delta S_{2}=(S_{1}\setminus S_{2})\cup(S_{2}\setminus S_{1})$ for their symmetric difference.
@@
Let $\mathcal{G}=(V,E)$ and $\mathcal{H}=(V^{\prime},E^{\prime})$ be finite graphs with $\mathcal{H}\subset\mathcal{G}$. For a current $\mathbf{n}$ on $\mathcal{G}$, let $\mathbf{n}_{|\mathcal{H}}$ denote its restriction to edges of $\mathcal{H}$. Then, for any $A\subset V$, $B\subset V^{\prime}$, $x,y\in V^{\prime}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}$,
$Z^{A,B}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}$ $=Z^{A\Delta\{x,y\},B\Delta\{x,y\}}_{\mathcal{G},\mathcal{H}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}.$ (3.6)
@@
$\sum_{\partial\mathbf{m}=A\Delta B}w(\mathbf{m})\mathds{1}\{\mathbf{m}\in\mathcal{E}\}\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.7)
We view $\mathbf{m},\mathbf{n}$ as multigraphs $\mathcal{M},\mathcal{N}$, and write $\partial\mathcal{N}\subset V^{\prime}$ for the vertices with an odd number of incident edges in $\mathcal{N}$. From this viewpoint, the sum over $\mathbf{n}$ above exactly counts the number of subgraphs $\mathcal{N}$ of $\mathcal{M}$ with $\partial\mathcal{N}=B$.
@@
$\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}=\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{H}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\Delta\{x,y\}\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.8)

### v2

@@ Abstract
We consider the Ising model on a $d$-dimensional discrete torus of volume $r^{d}$, in dimensions $d>4$ and for large $r$, in the vicinity of the infinite-volume critical point $\beta_{c}$. We prove that for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ (with a suitable constant) the susceptibility is bounded above and below by multiples of $r^{d/2}$. Additionally, again for $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$, the two-point function has a “plateau”: it decays like $|x|^{-(d-2)}$ when $|x|$ is small relative to the volume, but for larger $|x|$, it levels off to a constant value of order $r^{-d/2}$. We also prove that at $\beta=\beta_{c}-{\rm const}\,r^{-d/2}$ the renormalised coupling constant is nonzero, which implies a non-Gaussian limit for the average spin. The proof relies on near-critical estimates for the infinite-volume two-point function obtained recently by Duminil-Copin and Panis, and builds upon a strategy proposed by Papathanakos. The random current representation of the Ising model plays a central role in our analysis.
@@ 1.1 Introduction
In recent years, the plateau for PBC has been proven to exist for simple random walk in dimensions $d>2$ [43, 16], self-avoiding walk for $d>4$ [43, 30], (spread-out) percolation for $d>6$ [27, 28], and (spread-out) lattice trees and lattice animals for $d>8$ [32]. A theory of the effect of FBC vs PBC in the setting of the weakly-coupled hierarchical $n$-component $|\varphi|^{4}$ model in dimensions $d\geq 4$ was developed in [35, 40], and we comment further on this below. A general theory is outlined in [31].
For the Ising model in dimensions $d>4$, the absence of a plateau for FBC is an immediate consequence of (1.6) and the monotonicity of $\tau^{\Lambda_{r}}_{\beta_{c}}(0,x)$ in $r$. A matching lower bound for FBC is proved in [14]. For PBC, incomplete steps were taken in [39] towards proving existence of the plateau. In this paper, we build on the method of [39] to prove (1.8) for PBC in dimensions $d>4$. We also prove the non-vanishing of the renormalised coupling constant under PBC, indicating a non-Gaussian limit for the average spin. This can be contrasted with the Gaussian limit under FBC obtained in [2, 20, 4, 37].
We emphasise that we do not use the lace expansion in any way, and our results do not require a small parameter. The input from the lace expansion used to establish plateau theorems for self-avoiding walk, percolation, and lattice trees and lattice animals in [43, 30, 27, 28, 32] is replaced here by the near-critical bounds (1.6)–(1.7).
@@ 1.2 Main results
Let $d>4$. There are constants $c_{i}>0$, depending on $d$ but not on $\beta$ or $r$, such that
@@ 1.3 Conjectured asymptotic formulas
A precise and detailed account of the effect of boundary conditions on finite-size scaling is given in [35, 40], via a rigorous renormalisation group analysis. The results of [35, 40] are proved for the weakly-coupled $n$-component $|\varphi|^{4}$ model on the $d$-dimensional hierarchical lattice for $d\geq 4$ and for all $n\geq 1$, but, on the basis of universality, the results obtained there for the $1$-component model are conjectured to apply also to all $1$-component spin models in the universality class of the nearest-neighbour Ising model. An antecedent of the conjectures goes back to [13] (see also [46, Chapter 32]). The conjectures present predictions for a specific and universal non-Gaussian limit for the average field, and precise asymptotic amplitudes for the susceptibility, the plateau, and the renormalised coupling constant, throughout a critical window of width $r^{-d/2}$ (with a logarithmic correction for $d=4$) centred at $\beta_{c}$. We define the window scale and the large-field scale, respectively, by
@@
The function $f$ is strictly monotone increasing, and satisfies $f(s)\sim|s|^{-1}$ as $s\to-\infty$ and $f(s)\sim s$ as $s\to\infty$ (see [35, Section 1.5.2] where $f_{1}(s)=f(-s)$). The following conjecture proposes a precise non-Gaussian limit for the average field $S_{r}$, and a precise role for the universal profile $f$ in the critical finite-size scaling of the susceptibility, the two-point function, and the renormalised coupling constant for the Ising model. The conjecture is a special case ($n=1$) of conjectures from [35, 40] for $n$-component models for all $n\geq 0$ ($n=0$ corresponds to the self-avoiding walk). The profile $f$ also arises in the Curie–Weiss model; a recent example is [10, Theorem 1.1].
@@
As observed in [35, Section 1.5.3], the right-hand side of (1.20) is strictly monotone increasing, and has limits 00 and $2$ as $s\to-\infty$ and $s\to+\infty$ respectively. Thus, assuming (1.20), the renormalised coupling constant varies monotonically over its possible range $[0,2]$ as the window is traversed. At the infinite-volume critical point, (1.20) gives
@@
such that the statements of Conjecture 1.5 hold as stated when $\beta_{c}$ is replaced by $\beta_{c,r}$. In particular, we conjecture that the plateau phenomenon also occurs for free boundary conditions at the effective critical point.
@@ 1.4 Structure of proof
Let $d\geq 1$, $r\geq 3$, $\beta>0$, and $x\in{\mathbb{T}_{r}}$. Then $\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq T_{\beta}^{(r)}(x).$
@@
The upper bound follows from (1.9), $\tau_{\beta}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ by (1.6), and the fact that $\chi(\beta_{*})\asymp r^{d/2}$ from (1.5).
@@
Let $d\geq 1$. For all $r\geq 3$ and for all $\beta>0$,
$\sum_{x,y,z\in\mathbb{T}_{r}}|U_{4,\beta}^{\mathbb{T}_{r}}(0,x,y,z)|\geq\frac{1}{128d^{2}}\cdot\min\Bigl{\{}\beta(\chi^{\mathbb{T}_{r}})^{2}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta},\frac{1}{{\sf B}}\Big{(}\frac{\partial\chi^{\mathbb{T}_{r}}}{\partial\beta}\Big{)}^{2}\Big{\}}.$ (1.34)
@@ 2.2 Diagrammatic estimates
Let $d\geq 1$, $r\geq 3$, and $\beta>0$. For any $x\in{\mathbb{T}_{r}}$,
$T_{\beta}^{(r)}(x)-\tau_{\beta}^{\mathbb{T}_{r}}(x)\leq 2\Big{(}T_{\beta}^{(r)}\star\big{[}\big{(}T_{\beta}^{(r)}\big{)}^{2}[T_{\beta}^{(r)}-(\tau_{\beta}\circ\pi)]\big{]}\star T_{\beta}^{(r)}\Big{)}(x).$ (2.10)
@@
Since every point in ${\mathbb{Z}^{d}}$ can be written uniquely as a point in $\Lambda_{r}+r{\mathbb{Z}^{d}}$, we can write
@@
so the two sums over $y,z$ become four sums over $\check{y},\check{z},u,v$. Since $T_{\beta}^{(r)}$ is periodic, we have $T_{\beta}^{(r)}(y-z)=T_{\beta}^{(r)}(\check{y}-\check{z})$ and $T_{\beta}^{(r)}(x-y)=T_{\beta}^{(r)}(\check{x}-\check{y})$, independent of $u$. Hence, the sum over $u$ gives the upper bound
@@
It is interesting to compare the diagrammatic upper bound (2.10) with diagrams in related models. The counterpart of $T_{\beta}^{(r)}-\tau_{\beta}^{\mathbb{T}_{r}}$ is denoted by $\psi-\psi^{\mathbb{T}}$ in [43, 28, 30]. For weakly or strictly self-avoiding walk, the bounds [43, (6.36)] and [30, (4.20)] can be written as
@@
It follows from $\tau_{\beta}(x)\leq\tau_{\beta_{c}}(x)\lesssim|\!|\!|x|\!|\!|^{-(d-2)}$ and a standard convolution estimate [23, Proposition 1.7] that $(\tau_{\beta}*\tau_{\beta}^{2}*\tau_{\beta})(x)\lesssim|\!|\!|x|\!|\!|^{-(d-4)}$ when $d>4$, so this term can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ with $c_{5}$ large enough, uniformly in $\beta\leq\beta_{c}$. Also, for $\beta\leq\beta_{c}-c_{4}r^{-d/2}$, by (1.5) we have $r^{-d/2}\chi(\beta)\lesssim c_{4}^{-1}$, and we can make this as small as desired by taking $c_{4}$ large enough. With (2.18), this shows that the factor multiplying ${\chi(\beta)}/r^{d}$ in (2.17) can be made as small as desired by taking $\lVert x\rVert_{\infty}\geq c_{5}$ and $c_{4},c_{5}$ large. In particular, it can be made smaller than $\frac{1}{2}c_{T}$, which completes the proof. ∎
@@
In this section, we introduce a new coupling between Ising models on ${\mathbb{T}_{r}}$ and on large finite subsets of $\mathbb{Z}^{d}$. The coupling uses the random current representation (see [17] or [38, Chapter 2] for an introduction) and a percolation exploration process. To derive useful bounds, we develop a new switching lemma between currents on $\mathbb{T}_{r}$ and on large finite subsets of $\mathbb{Z}^{d}$, in Lemma 3.4.
@@ 3.1 Random currents and switching lemma
$w_{\beta}(\mathbf{n})=\prod_{e\in E}\dfrac{\beta^{\mathbf{n}_{e}}}{\mathbf{n}_{e}!},\qquad Z^{A}_{\mathcal{G},\beta}=\sum_{\partial\mathbf{n}=A}w_{\beta}(\mathbf{n}).$ (3.3)
@@
Observe that, by definition, $Z_{\mathcal{G},\beta}^{A}[\Omega_{\mathcal{G}}]=Z^{A}_{\mathcal{G},\beta}$.
@@
Given two graphs $\mathcal{G},\mathcal{H}$, two vertex sets $A\subset\mathcal{G}$, $B\subset\mathcal{H}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}\times\Omega_{\mathcal{H}}$, we write
@@
We often drop the subscript $\beta$ when it does not play a role. For two sets $S_{1},S_{2}$, we write $S_{1}\Delta S_{2}=(S_{1}\setminus S_{2})\cup(S_{2}\setminus S_{1})$ for their symmetric difference. For two currents $\mathbf{n}$ and $\mathbf{m}$, we write $\mathbf{n}\leq\mathbf{m}$ when $\mathbf{n}_{e}\leq\mathbf{m}_{e}$ for each edge $e$, and we write $\binom{\mathbf{m}}{\mathbf{n}}=\prod_{e\in E}\binom{\mathbf{m}_{e}}{\mathbf{n}_{e}}$ for the product of the binomial coefficients. We will use the following minor extension of the switching lemma of [21, 2] (see also [17, Lemma 7]).
@@
Let $\mathcal{G}=(V,E)$ and $\mathcal{H}=(V^{\prime},E^{\prime})$ be finite graphs with $\mathcal{H}\subset\mathcal{G}$. For a current $\mathbf{n}$ on $\mathcal{G}$, let $\mathbf{n}_{|\mathcal{H}}$ denote its restriction to edges of $\mathcal{H}$. Then, for any $A,B\subset V$, $x,y\in V^{\prime}$, and $\mathcal{E}\subset\Omega_{\mathcal{G}}$,
$Z^{A,B}_{\mathcal{G},\mathcal{G}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}$ $=Z^{A\Delta\{x,y\},B\Delta\{x,y\}}_{\mathcal{G},\mathcal{G}}\Big{[}\mathbf{n}_{1}+\mathbf{n}_{2}\in\mathcal{E},\>x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}.$ (3.6)
It is an immediate consequence of (3.6) that
$Z^{A,B}_{\mathcal{G},\mathcal{G}}\Big{[}x\xleftrightarrow{(\mathbf{n}_{1}+\mathbf{n}_{2})_{|\mathcal{H}}\>}y\Big{]}\leq Z^{A\Delta\{x,y\}}_{\mathcal{G}}Z^{B\Delta\{x,y\}}_{\mathcal{G}}.$ (3.7)
@@
$\sum_{\partial\mathbf{m}=A\Delta B}w(\mathbf{m})\mathds{1}\{\mathbf{m}\in\mathcal{E}\}\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.8)
We view $\mathbf{m},\mathbf{n}$ as multigraphs $\mathcal{M},\mathcal{N}$, and write $\partial\mathcal{N}\subset V$ for the vertices with an odd number of incident edges in $\mathcal{N}$. From this viewpoint, the sum over $\mathbf{n}$ above exactly counts the number of subgraphs $\mathcal{N}$ of $\mathcal{M}$ with $\partial\mathcal{N}=B$.
@@
$\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}=\mathds{1}\{x\xleftrightarrow{\mathbf{m}_{|\mathcal{H}}\>}y\}\sum_{\begin{subarray}{c}\mathbf{n}\in\Omega_{\mathcal{G}},\>\mathbf{n}\leq\mathbf{m}\\ \partial\mathbf{n}=B\Delta\{x,y\}\end{subarray}}\binom{\mathbf{m}}{\mathbf{n}}.$ (3.9)